A Bibliography of Publications of James H. Wilkinson

Sven Hammarling The Numerical Algorithms Group

Wilkinson House Jordan Hill Road Oxford OX2 8DR

UK

Tel: ?n/a? FAX: ?n/a?

E-mail: na.hammarling@na-net.ornl.gov (Internet)

08 August 2024 Version 1.83

Abstract

This bibliography records publications of the late James H. Wilkinson.

Title word cross-reference

$3.70 [I.70]. A [PW69]. Au = λBu [PW74c, Wil77a]. Ax = λBx [GUW70, GUW72, MW68a, MW71b, PW69, PW70a, Wil76]. B [PW69]. Bẋ = Ax [Wil77a]. I − 2ww^{H} [Wil65e]. LR [MW68d, MW71a, PW70b, PW71b, Wil65d]. QL [BMRW68, BMRW71, DMW71, MW68c]. QR [BMRW68, BMRW71, MRW70, MPW70, MRW71a, MPW71a, PW70b, PW71b, Wil65a, Wil65d, Wil68c, Wil69b]. QZ [Wil78b, Wil79].

100s [Wil67c]. 12th [Jac76]. 13-15 [Cra87]. 15th [Jac76]. 16-19 [MTTT63]. 1954 [Wil55c]. 1962 [MTTT63]. 1977 [Pow77]. 1986 [Fox87a, IP87]. 1987 [Cra87]. 19th [Jac78].

2 [Wil61e]. 22nd [Jac78]. 25 [DHW90].

6th [Pow77].

90s [Wil70c].

Accuracy [DMW81, DMW83, PW74a]. ACE [Wil75c, Ano12, Cop12, Wil51b, Wil52, Wil54a, Wil54b, Wil54c, Wil55a, Wil80b, Wil05, Wil12]. ACM [Cra87, Ash87]. Admissible [OPW64, OPW65]. Advances [Wil77b, Wil80a, dBG78, HHT19]. Alan [Ano12, CK12, Cop05, Cop12]. Algebra [HHT19, Wil54b, WR71, Wil74c, Wil74e, Wil77b, Wil78e, Wil86b, Wil80a, Par72]. Algebraic [FHW48b, Gra66, Par66, Wil66a, Wil60b, Wil60d, Wil63d, Wil65b, Wil67e, Wil68d, Wil70a, Wil71c, Wil88, Wil23, FHW48a, Wil67b, Wil94]. Algebraicheskaya [Wil70a]. Algebraicznych [Wil67b]. Algébraiques [Wil62a, Wil63a]. Algorithm [GUW70, GUW72, MW68c, MW68d, MPW70, PW74b, Wat11, Wil65a, Wil68c, Wil69b, Wil78b, Wil79, DMW71, MRW70, MRW71a, MPW71a, MW71a]. Algorithms [BMRW68, PW74c, Wil67f, Wil68e, Wil74c, Wil74g, Bar02, BMRW71, Wil65d, Wil61e, Wil70b]. Almost [SW81b, Wil58d, Wil59c, Wil67a, Wil68a]. Alston [Wil75a, Wil78a]. American [MTTT63]. Analyser [Wil51a]. Analysis [Gol84, IP87, Par90, Wil60a, Wil61a, Wil61b, Wil62d, Wil65e, Wil67c, Wil68b, Wil68d, Wil68e, Wil71a, Wil71e, Wil72c, Wil74a, Wil84c, Wil85, Wil86a, Jac76, Par87, Pow77, Sca74, dBG78]. Analyst [Wil71f, Wil87c]. Applications [Jac78, JRW62]. Applied [MTTT63]. Approximations [HW80]. April [IP87, Jac76, MTTT63]. Arising [HW80, PW71c]. Arithmetic [Wil63c, MTTT63, RR23]. Art [IP87, Wil84c, Wil85, Jac76]. Aspect [Wil57h]. Aspects [Wil78f]. Assessment [Wil55a]. Associated [SW80a, SW80b]. Atlantic [MTTT63]. August [ICM75]. Automatic [Par72, Cop05, DW59, Wil48a, Wil48b, Wil51c]. availability [Jac78]. Award [Ash87, Wil75a, Wil78a].

B. [TW12]. Band [MW65, MW67, PW69, MRW70, MRW71a, MW71e, MW71d]. Based [Wil62d, Wil65e, Wil67f]. Basic [Wil68b, Wil78f]. Behaviour [HW76a]. Birmingham [IP87]. Birth [HH19a, NW85]. Bisection [BW67, BMW71, Wil62b]. Bishop [Wil67c]. Bledy [Wil67b]. Book [Gra66, I.70, Par66, Par72, Wil66a, Wil62a, Wil63a, Wil67c, Wil70b, Wil70c]. bound [Bar03]. Bounds [OW82, SW80a, SW81a, Wil61f, Bar02, SW80b]. Brain [Cop12].

Britain [Wil57c]. British [Wil57a, WCK89]. Broyden [PW74b]. Build [Cop12, Cop05]. Bulirsch [Mol22].

C [Par72]. Calculation

[BW67, BMW71, MW67, Wil54a, Wil57b, Wil58b, Wil58a, Wil62b, Wil62c, Wil65c, Wil66b, MW71d, PW71a]. Cambridge [Wil67c]. Canada [ICM75]. Canonical [GW75, GW76, Wil78b, Wil78c, Wil79]. Card [FHW48b]. Celebrating [HH19a, HH19c]. Centenary [HH19a]. Center [dBG78]. Century [MHR80]. Chicago [MTTT63]. Christian [Mol22]. Cie [Wil62a, Wil63a]. City [MTTT63]. Classical [Wil74a]. Clenshaw [Bar02]. Cleve [Mol22]. Close [Wil67a, Wil68a]. cm [I.70]. co [RW61]. co-diagonalization [RW61]. codebreaker [Cop05]. Codiagonal [Wil57b, Wil58b]. Coding [Wil51b, Wil52, Wil55a]. Coefficients [OPW64, OPW65]. Collection [MHR80, Wil70b]. Comments [Wil71f, Wil87c]. Complex [BMPW66, MW68d, PW70b, BMPW71, MW71a, PW71b]. Computation [Cra87, GW75, GW76, Par72, Ral65a, Ral65b, Wil60a, CH90]. Computational [Wil70b]. Computed [DMW81, DMW83, PW74a, SW81a, Wil61f]. Computer [BN71, Cop12, NW85, Wil57a, WCK89, Cop05, RW61]. Computers [JRW62, Wil74e]. Computing [CGM^{+}61, Cop05, DW59, DHW90, FGM^{+}57, FGM^{+}70, MW81, MHR80, Nas90, Wil48a, Wil48b, Wil51c, Wil57c, DHW92, HNW^{+}48, MTTT63, Wil55b]. Condition [CMSW79, Wil86b]. Conditioned [GW75, GW76, PW79, Wil67d, Wil74d, Wil59a, Wil59b, Wil72b]. conducted [dBG78]. Conference [Cra87, HW76b, IP87, Jac76, Jac78, MW81, HHT19]. Conferences [WCK89]. Congress [ICM75]. Considerations [DHW90, DHW92].

Construction [Wil80b]. Contribution [Par90, Pow77, Par87]. Convergence [Wil65d, Wil68c, Wil69b, Wil62g]. Copeland [TW12]. Cornelius [Sca74]. Corner [Mol22]. Coupling [DGKW84b]. cyclic [Wil62g].

D [Wil67c, Wil70b]. Data [Wil51a]. December [Wil55c]. Decomposition [MW65, MPW65, Wil74f, Wil78f, MPW71c, MW71e, RR23]. Defective [SW81b]. Defined [OPW64, OPW65]. Definite [MW65, MPW65, MPW66, MPW71b, MPW71c, MW71e]. Deflation [PW74b]. Design [Wil51c, Wil57a]. DEUCE [Wil57d, Wil80b]. Diagonal [Wil62f, Wil67a, Wil68a]. diagonalization [RW61]. Difference [HW80]. Differential [HW80, Wil57e, Wil61c, Wil71a, Wil76, Wil77a, Wil78c]. Digital [JRW62, Ral65a, Ral65b, Wil57a, Wil74e]. Direct [Wil61b, Wil67e]. Discussion [Wil71a, HNW^{+}48, Wil69b]. divisors [Wil84a]. double [RR23]. Dr. [Fox87b, Pow77]. Drazin [Wil78d, Wil82]. Durand [Wil62a, Wil63a].

Early [WCK89]. edited [TW12]. Editor [RW62]. efficient [RW61]. Eigenpairs [FP20]. Eigenproblem [MW68a, PW70a, Wil60b, MW71b, Wil72b]. Eigensystems [GW75, GW76, PW74a, Wil61f, Wil66b, Wil74d]. Eigenvalue [DGW82, DGKW84a, Par66, PW74c, SW80a, Wil66a, Wil62d, Wil65b, Wil70a, Wil71c, Wil77a, Wil87a, Wil88, SW80b]. Eigenvalues [BW67, DMW81, DMW83, HVW70, PW69, Wil65a, Wil67a, Wil68a, Wil84b, Wil86c, BMW71, Wil62b]. Eigenvector [SW80a, SW80b]. Eigenvectors [DMW81, DMW83, MW67, PW70b, PW71b, Wil57b, Wil58b, Wil58a, MW71d, PW71a, Wil62c]. Electronic [Cop12, Wil55b]. Elementary [Wil58d, Wil59c, Wil70d, Wil84a]. Elimination [OW82, PW75, Wil62f]. Elliptic [HW80]. Engine [Wil51c, Cop05, DW59, Wil48a, Wil48b]. Engineering [JRW62]. English [I.70]. Equation [DGKW84b]. Equations [BMPW66, FHW48b, HW80, MPW66, MW67, PW71c, PW79, Wil57e, Wil57g, Wil60c, Wil61c, Wil61d, Wil67d, Wil67e, Wil71a, Wil76, Wil78c, BMPW71, FHW48a, MPW71b, MW71d, Wil70c, Wil62a, Wil63a]. Error [OW82, Ral65a, Ral65b, SW80a, SW81a, Wil60a, Wil61a, Wil61b, Wil61f, Wil62d, Wil65e, Wil68b, Wil68d, Wil68e, Wil71e, Wil74a, Wil84c, Wil85, Wil86a, Bar02, Bar03, SW80b]. Errors [Gra66, Wil60d, Wil63b, Wil63d, Wil23, Wil67b, Wil94]. Essays [MHR80]. Estimate [CMSW79]. evaluation [Bar02, Bar03, Wil59a, Wil59b]. Example [CS20, Wil57d]. Examples [BN71, CS20]. Expanded [Wil74b]. Experimental [MTTT63]. Expounded [Wil74b]. extended [RR23].

F.R.S. [Fox87b]. Facilities [Wil57c]. Fastest [Cop12]. Feingold [Wil87b]. fifteenth [MTTT63]. Fifth [HW76b]. Finding [Wil55d]. Finite [HW80]. first [Ash87]. Floating [Wil60a, Wil63c]. Floating-Point [Wil60a, Wil63c]. Ford [PW74b]. Foreward [FW75]. Form [GW75, GW76, MW68a, MW68b, Wil62f, Wil78b, Wil78c, Wil79, MW71b, MW71c, Wil65e]. Forms [Wil58d, Wil59c, Wil68b]. Forsythe [Bar02]. Francis [Wat11].

G [I.70, Wil67c]. Gauss [PW75]. Gaussian [OW82]. General [MW68b, MW71c]. Generalization [Wil70d]. Generalized [PW70a, PW74c, Wil77a]. German [I.70]. Gerschgorin [HVW70]. Givens [RW61, RW62, Wil57b]. Gladwell [Wil67c]. Global [Wil68c, Wil69b]. Goos [I.70]. GPO [Wil51a]. Great [Wil57c]. Gregory [Wil70b].

H [CDD87, Eva76, Fox87b, Gra66, HH19a, Par66, Par87, Par90, Par72, Pow77]. Handbook [Wil70c, Wil74b, Par72]. Hardy [Fox87a, HH19c, Cle04]. held [IP87, Jac76, Jac78, MTTT63, Pow77]. Hermitian [DGW82, DGKW84a]. Hessenberg [MW68d, MW68b, MPW70, MPW71a, MW71a, MW71c]. high [MTTT63]. Highlights [HHT19]. History [Cra87, MHR80, Nas90]. Hoffman [Wil70d]. honour [Sca74]. Householder [Wil78a, MRW68, MRW71b, RW62, Wil60b, Wil62e, Wil75a]. HSNC’87 [Cra87]. Hyperbolic [Wil57e, Wil61c].

II [Par72, Wil63a, Wil59b, Wil71c, Wil86c]. II. [Wil51c]. Ill [GW75, GW76, PW79, Wil67d, Wil74d, Wil86b, Wil59a, Wil59b, Wil72b]. Ill-Condition [Wil86b]. Ill-Conditioned [GW75, GW76, PW79, Wil67d, Wil74d, Wil59a, Wil59b, Wil72b]. Illinois [MTTT63]. IMA [IP87]. IMA/SIAM [IP87]. Implicit [MW68c, DMW71]. Improving [DMW81, DMW83]. Including [DGKW84b]. Inclusion [HVW70]. Infinite [Wil57h]. Instability [Wil62f]. International [ICM75]. Interpolation [CS20, Wil67f]. Interview [Eva76]. Introduction [Wil71c, Wil71b, TW12]. Introductory [Wil71d, Wil71a]. Invariant [DHW90, PW74a, SW81a, Wil75b, DHW92]. Inverse [PW79, Wil72a, Wil74d, Wil78d, Wil82, PW71a, Wil62c]. Inverses [PW70c]. Inversion [DGKW84b, Wil61b, Wil70c]. Inverting [Wil74g]. Iteration [PW79, Wil72a, Wil74d, PW71a, Wil62c]. Iterative [GW66a, GW66b, HW76a, MPW66, MPW71b, Wil55d].

J [Fox87b, Gra66, Par66, Par72, Pow77]. J.

[Eva76, I.70, Par87, Par90]. Jack [TW12]. Jacobi [Wil62g]. James [Cle04, CDD87, Fox87a, HH19a, HH19c]. Jersey [Cra87, MTTT63]. Joan [Wil70c]. John [Wil70b, Wil70c]. Joint [IP87]. Jordan [GW75, GW76, PW75]. July [Pow77].

Karney [Wil70b]. Kronecker

[Wil78b, Wil78c, Wil79].

L [Wil67c, Wil70b]. Laboratory

[DW59, Wil48a, Wil55a, Wil58c, Wil72c, Wil80b, Wil05, Wil12]. Lamé [Wil65c]. Lanczos [GUW70, GUW72, Sca74, Wil57b, Wil58a]. LAPACK [DHW90]. Large [Wil63b]. Large-Scale [Wil63b]. Latent [Wil54a, Wil55d, Wil57f]. Least [GW66a, GW66b, PW70c, Wil71b]. lecture [TW05, TW12]. Lectures [Ash87]. Letter [RW62]. level [RW61]. Linear [BMPW66, FHW48b, HW76a, HW80, HHT19, OPW64, OPW65, Par72, Wil54b, Wil57g, Wil60c, Wil61d, Wil67d, Wil67e, Wil67f, Wil70c, Wil71b, WR71, Wil74a, Wil74c, Wil74e, Wil74g, Wil76, Wil77b, Wil77c, Wil78c, Wil78e, Wil86b, BMPW71, FHW48a, Wil80a]. Loading [Wil57h]. Logical [Wil51c]. London [Wil67c, Pow77]. Ltd [Wil70b].

M [Wil67c]. machine [CK12]. machines [HNW^{+}48, Wil55b]. Madison [dBG78]. Manitoba [HW76b, MW81]. Masson [Wil62a, Wil63a]. master [Cop05]. Mathematical [MTTT63]. mathematician [Cle04]. Mathematicians [ICM75]. Mathematics [Eva74, HW76b, MW81, dBG78, MTTT63]. Matrices [BMRW68, DGW82, DGKW84a, FP20, HVW70, MW65, MW67, MW68d, PW70b, SW81b, Wil54a, Wil55d, Wil57b, Wil57g, Wil57f, Wil58b, Wil61d, Wil65a, Wil65e, Wil66b, Wil67a, Wil68a, Wil70b, Wil74g, BMRW71, MRW70, MPW70, MRW71a, MPW71a, MW71a, MW71e, MW71d, PW71b, Wil62e, Wil72b, Wil84a]. Matrix [CMSW79, MPW65, MW68b, Wil58c, Wil58d, Wil59c, Wil61b, Wil62f, Wil67c, Wil67e, Wil68b, Wil68e, Wil70c, BW67, BMW71, MRW68, MRW71b, MPW71c, MW71c, RW61, Wil62b, Wil62c]. May [Cra87, dBG78]. Memory [CDD87]. Method [BW67, PW79, Wil58a, Wil60b, Wil62f, BMW71, RW61, Wil62b, Wil62e]. Methods [CGM^{+}61, FHW48b, FGM^{+}57, FGM^{+}70, HW76a, Wil55d, Wil60c, Wil61b, Wil67e]. Michaelson [Wil67c]. minfit [Wil74f]. Model [Wil51b, Wil51c, Wil52, Wil54a]. Modern [CGM^{+}61, FGM^{+}57, FGM^{+}70, Wil71e, Cop05]. Modified [MW68d, MW71a]. Multiple [Wil65a, Wil67a, Wil68a].

National [DW59, Wil48a, Wil55a, Wil58c, Wil72c, Wil80b, Wil05, Wil12]. needs [Jac78]. neighbouring [Wil84a]. Newton [PW79]. NF [Wil62a, Wil63a]. No [DHW90]. Note [DHW90, GW66b, HVW70, Wil62g, Wil72b, Wil74d, Wil78d, Wil82]. Notes [FHW48a]. NPL [Wil75c]. Number [CMSW79]. Numeric [Cra87]. Numerical [DHW90, DHW92, Eva74, Gol84, HHT19, HW76b, IP87, Jac78, MW81, Par90, Wil63b, Wil70c, Wil71a, Wil71f, Wil72c, Wil74e, Wil77b, Wil78e, Wil86b, CH90, Jac76, Par87, Pow77, Sca74, Wil80a, Wil87c, dBG78]. Numériques [Wil62a, Wil63a].

Obituary [Fox87b]. October [Fox87a, HW76b, MW81]. Optimum [Wil55a]. Orbit [DGKW84b]. Origin [Wil68c]. Orthogonal [Wil62d, Bar02]. Oscillating [Wil57h]. other [CK12].

p139 [Wil74b]. pages [Wil67c, Wil70b]. paperbound [I.70]. papers [Cra87, Sca74]. Paris [Wil62a, Wil63a]. Part [Wil71c, Wil71b]. Partial [Wil57e, Wil61c, Wil71a]. Particular [HW76a]. Perfidious [Wil84d]. Physical [DW59, Wil48a, Wil55a, Wil58c, Wil72c, Wil80b, Wil05, Wil12]. Pilot [Wil51b, Wil51c, Wil52, Wil54a, Wil54b, Wil54c, Wil55a, Wil75c, Wil80b, Wil05, Wil12]. Pivoting [PW75]. Plane [Wil63c]. Point [Wil60a, Wil63c]. Polish [Wil67b]. Polynomial [PW71c, PW74b, Wil84d, Bar02, Bar03]. Polynomials [Wil65c, Wil59a, Wil59b]. Positive [MW65, MPW65, MPW66, MPW71b, MPW71c, MW71e]. Posteriori [OW82]. pp [I.70, Wil62a, Wil63a, Wil70c]. Practical [HW76a, PW71c, Wil78d, Wil82]. Practice [Wil72a]. Precision [Wil77c, RR23]. presented [Cra87]. Press [Wil67c]. Price [I.70]. Princeton [Cra87]. Priori [Wil68d]. Problem [DGW82, DGKW84a, GUW72, PW70c, PW74c, Wil65b, Wil70a, Wil71c, Wil77a, Wil78e, Wil88, GUW70, Par66, Wil66a]. problema [Wil70a]. Problems [MW68a, PW71c, Wil63b, Wil67e, Wil87a, MW71b]. Proceedings [HW76b, ICM75, IP87, Jac76, MW81, MTTT63, Pow77, Cra87, Jac78, dBG78]. Procesach [Wil67b]. process [Wil62g]. Processes [Wil57b, Wil60d, Wil63d, Wil67b, Wil68b, Wil68d, Wil23, RW62, Wil94, Gra66]. Produced [Wil57b]. Programmes [Wil51a]. Programming [Wil51b, Wil52, Wil71b]. Programs [Wil58c]. Progress [Wil48b, Wil57a]. Proof [Wil70d]. Pseudo [PW70c]. Pseudo-Inverses [PW70c]. Punched [FHW48b].

quadratic [Wil62g, Wil84a]. Quotes [HH19b].

R [Wil67c]. Radar [Wil51a]. Ratio [Wil57h]. Readings [BN71]. Real [BMPW66, PW70b, Wil65a, BMPW71, MPW70, MPW71a, PW71b]. Realistic [SW80a, SW80b]. Reducing [Wil62f]. Reduction [MW68a, MW68b, MW71b, Wil58d, Wil59c, MW71c]. Reference [HW76a]. Refinement [GW66a, GW66b, MPW66, MPW71b]. Reinsch [Par72, Mol22]. Related [MW68a, MW71b, Wil65d]. Reliable [CH90]. Remarks [Wil61e, Wil71a, Wil71d]. Report [Wil48b, Wil51c, Wil55c]. Research [dBG78]. Residual [Wil77c]. Reversal [DGW82, DGKW84a, DGKW84b]. Review [Gra66, I.70, Par66, Par72, Wil66a, Wil62a, Wil63a, Wil67c, Wil70b, Wil70c]. Revisited [Wil86a]. Rigorous [Wil61f]. Robert [Wil70b]. Roland [Mol22]. Rootfinding [CS20]. Roots [Wil54a, Wil55d, Wil57f]. Rotations [Wil63c]. Rounding [Bar02, Wil60d, Wil63d, Wil94, Wil23, Bar03, Wil67b, Gra66]. Rundungsfehler [I.70, Wil69a]. Runge [CS20]. Russian [Wil70a].

S [Wil67c, Wil75a, Wil78a]. S.O.R. [HW76a]. Scale [Wil63b]. scheme [RW61]. Scientific [Cra87, JRW62, Nas90]. Secular [DGKW84b]. Sensitivity [Wil84b, Wil86c]. September [Fox87a, Jac78, Wil55c]. series [Bar02, TW05, TW12]. Sharply [OPW64, OPW65]. Shifts [Wil68c]. SIAM [IP87]. Significance [Wil78d, Wil82]. Similarity [MW68b, MW71c, Wil58d, Wil59c]. Simple [SW80a, SW80b]. Simultaneous [FHW48b, FHW48a]. Single [Wil77c]. Single-Precision [Wil77c]. Singular [RR23, Wil74f, Wil78f]. Singular-Value [Wil78f]. sobstvennyh [Wil70a]. Society [MTTT63]. Software [Eva74, Jac78]. Solution [BMPW66, BMPW71, FHW48b, GW66a, GW66b, MPW66, MW67, MW71d, PW71c, PW74c, Wil60b, Wil67d, Wil67e, Wil70c, Wil74a, Wil77c, FHW48a, MPW71b]. Solutions [OPW64, OPW65, Wil62a, Wil63a]. Solving [DGKW84b, Wil60c, Wil74g]. Some [PW74c, Wil71f, Wil77b, Wil80a, Wil87c]. Sons [Wil70b, Wil70c]. Special [CDD87]. specified [PW71a]. speed [MTTT63]. Spin [DGKW84b]. Springer [I.70]. Springer-Verlag [I.70]. Square [GW66a]. Squares [GW66b, PW70c, Wil71b]. Stability [PW75, Wil58d, Wil59c]. Standard [MW68a, MW71b]. State [Wil84c, Wil85, Jac76, IP87]. store [RW61]. Stream [Wil57h]. Structures [BN71]. Struggle [Cop12, Cop05]. Studies [Gol84, Sca74]. Subsonic [Wil57h]. Subspaces [DHW90, PW74a, SW81a, Wil75b, DHW92]. Successive [Wil67f]. Survey [Wil68e]. Sussex [Jac78]. SVD [Mol22]. Symmetric [BW67, BMRW68, GUW70, GUW72, MW65, MPW65, MW67, MW68a, MPW71c, MW71e, PW69, PW74c, Wil65a, BMW71, BMRW71, MRW68, MRW70, MRW71a, MRW71b, MW71b, MW71d, RW61, Wil62b, Wil62c, Wil62e]. Symmetry [DGW82, DGKW84a, DGKW84b]. Symposium [MTTT63, Pow77, dBG78]. System [MPW66, Wil55a, Wil77a, MPW71b]. Systems [BMPW66, DGKW84b, HW80, OPW64, OPW65, Wil71b, Wil74a, Wil74g, Wil77c, BMPW71].

T [Wil70b]. Techniques [Wil62d]. tenth [MW81]. Test [Wil74g]. Testing [Wil70b]. their [Wil55b]. Theorem [HVW70, Wil70d, Wil87b]. Theory [Wil72a]. Time [DGW82, DGKW84a, DGKW84b]. Time-Reversal [DGKW84a, DGKW84b]. Transformations [Wil58d, Wil59c, Wil62d, Wil65e]. translated [I.70]. Tri [Wil62f]. Tri-Diagonal [Wil62f]. Triangular [Wil58d, Wil59c]. Triangularization [PW70b]. triangularizations [PW71b]. Tridiagonal [BW67, Wil68c, BMW71, Wil62b, Wil62c]. tridiagonalization [MRW68, MRW71b]. Turing [Ano12, Ash87, CK12, Cop05, Cop12, TW05, TW12, Wil80b]. Twentieth [MHR80]. twenty [Ash87]. Two [Wil67f, RW61]. two-level [RW61].

U.S.A. [Wil55c]. unified [Bar03]. universal [CK12]. University [IP87, Jac76, Jac78, Wil67c, dBG78]. unknown [Wil62h]. Unsymmetric [MW67, MW71d]. Use [Wil55d, Wil57d, Wil65e, Wil77c]. Used [Wil55a]. uses [Wil55b].

Value [Wil74f, Wil78f, RR23]. Vancouver [ICM75]. Vectors [Wil54a, Wil55d]. Verlag [I.70]. very [Wil72b]. Vibration [Wil67c]. Visit [Wil55c]. Vol [Par72]. Volume [CDD87, Wil63a].

w [Wil67b]. Westlake [Wil70c]. Wielandt [Wil70d]. Wiley [Wil70b, Wil70c]. Wilkinson [Fox87a, Fox87b, Gra66, I.70, Par66, Par72, CDD87, Cle04, CS20, Eva76, FP20, HH19a, HH19c, HH19b, Par87, Par90, Pow77, TW05, TW12]. Wing [Wil57h]. Wisconsin [dBG78]. Work [Wil80b]. Working [DHW90]. World [Cop12].

x [I.70].

years [Ash87]. York [I.70, Jac76, Wil70c].

Zaokragleń [Wil67b]. zeros [Wil59a, Wil59b]. znachenii [Wil70a].

References

Anonymous:2012:ATA

[Ano12]  Anonymous.  Alan Turing’s

ACE. Web site, May 28, 2012.

                                                   URL https://i-programmer.
                                                   info/history/machines/11-
                                                   an-ace-of-a-machine.html.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966– 1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Barrio:2002:REB

[Bar02] Roberto Barrio. Rounding error bounds for the Clenshaw and Forsythe algorithms for the evaluation of orthogonal polynomial series. Journal of Computational and Applied Mathematics, 138 (2):185–204, January 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/S037704270100382X.

Barrio:2003:URE

[Bar03] R. Barrio. A unified rounding error bound for polynomial evaluation. Advances in Computational Mathematics, 19 (4):385–399, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1024203520270.

Bowdler:1966:SRC

[BMPW66] H. J. Bowdler, R. S. Martin, G. Peters, and J. H. Wilkinson. Solution of real and complex systems of linear equations. Numerische Mathematik, 8(3):217–234, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See also [WR71, pp. 93–110].

Bowdler:1971:SRC

[BMPW71] H. J. Bowdler, R. S. Martin, G. Peters, and J. H. Wilkinson. Solution of real and complex systems of linear equations. In Wilkinson and Reinsch [WR71], pages 93–110. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Bowdler:1968:ASM

[BMRW68] H. J. Bowdler, R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR and QL algorithms for symmetric matrices. Numerische Mathematik, 11(4):293–306, 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See also [WR71, pp. 227–240].

Bowdler:1971:ASM

[BMRW71] H. Bowdler, R. S. Martin, C. H.

Reinsch, and J. H. Wilkinson. The QR and QL algorithms for symmetric matrices. In Wilkinson and Reinsch [WR71], pages 227–240. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Barth:1971:CES

[BMW71] W. Barth, R. S. Martin, and J. H. Wilkinson. Calculation of the eigenvalues of a symmetric tridiagonal matrix by the method of bisection. In Wilkinson and Reinsch [WR71], pages 249–256. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell, editors. Computer Structures: Readings and Examples. McGraw-Hill, New York, 1971. ISBN 0-07-004357-4. xix + 668 pp. LCCN TK7888.3 .B4.

Barth:1967:CES

[BW67] R. S. Martin W. Barth and J. H. Wilkinson. Calculation of the eigenvalues of a symmetric tridiagonal matrix by the method of bisection. Numerische Mathematik, 9(5):386–393, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 249– 256].

Chatelin:1987:SVM

[CDD87] F. Chatelin, J. Dongarra, and I. Duff. Special volume in memory of James H. Wilkinson. Linear Algebra and its Applications, 88–89(??):??, ???? 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
journal/00243795/88-89.

Clenshaw:1961:MCM

[CGM^{+}61] C. W. Clenshaw, E. T. Goodwin, D. W. Martin, G. F. Miller, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of National Physical Laboratory. Notes on Applied Science. HMSO, London, UK, second edition, 1961. vi + 170 pp. LCCN QA297 .T4 1961.

Cox:1990:RNC

[CH90] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Campbell-Kelly:2012:ATO

[CK12] Martin Campbell-Kelly. Alan Turing’s other universal machine. Communications of the ACM, 55(7):31–33, July 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clenshaw:2004:WJH

[Cle04]  C. W. Clenshaw.  Wilkin-

son, James Hardy (1919–1986), mathematician. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, September 2004.

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W. Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM Journal on Numerical Analysis, 16 (2):368–375, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Copeland:2005:ATA

[Cop05] B. Jack Copeland, editor. Alan Turing’s Automatic Computing Engine: the master codebreaker’s struggle to build the modern computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856593-3 (hardcover). xx + 553 pp. LCCN QA75 .A43 2005.

URL http://ukcatalogue.
oup.com/product/9780198565932.
do;   http://www.oxfordscholarship.
com/oso/public/content/maths/
9780198565932/toc.html.

Copeland:2012:ATE

[Cop12] B. Jack Copeland, editor. Alan Turing’s Electronic Brain: the Struggle to Build the ACE, the World’s Fastest Computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-960915-2 (paperback). xxi + 553 pp. LCCN ???? UK £14.99.

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87:

ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Corless:2020:REI

[CS20] Robert M. Corless and Leili Rafiee Sevyeri. The Runge example for interpolation and Wilkinson’s examples for rootfinding. SIAM Review, 62(1):231–243, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBoor:1978:RAN

[dBG78] Carl de Boor and Gene H.

Golub, editors. Recent advances in numerical analysis: proceedings of a symposium conducted by the Mathematics Research Center, the University of Wisconsin–Madison, May 22– 24, 1978, volume 41 of Publ. Math. Res. Center Univ. Wisconsin. Academic Press, New York, NY, USA, 1978. ISBN 0-12-208360-1. LCCN QA297 S994 1978.

Dongarra:1984:EPH

[DGKW84a] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. The eigenvalue problem for Hermitian matrices with time-reversal symmetry. Linear Algebra and its Applications, 60:27–42, 1984. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dongarra:1984:SSE

[DGKW84b] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. Solving the secular equation including spin orbit coupling for systems with inversion and time-reversal symmetry. Journal of Computational Physics, 54(2):278–288, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.netlib.org/utk/people/
JackDongarra/PAPERS/Inversion-
and-Time-Reversal-Symmetry.
pdf;   http://www.sciencedirect.
com/science/article/pii/0021999184901190.

Dongarra:1982:EPH

[DGW82] J. J. Dongarra, J. R. Gabriel, and J. H. Wilkinson. The eigenvalue problem for Hermitian matrices with time reversal symmetry. Technical Memorandum 3, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, Illinois 60439, USA, 1982. Published as [DGKW84a] where D. D. Koelling is also an author.

Dongarra:1990:NCC

[DHW90] J. J. Dongarra, S. Hammarling, and J. H. Wilkinson. Numerical considerations in computing invariant subspaces. LAPACK Working Note No. 25. Technical Report CS-90-117, Department of Computer Science, University of Tennessee, 107 Ayres Hall, Knoxville, TN 37996-1301, USA, 1990. Also produced as Oak Ridge National Laboratory Report, ORNL/TM-11704. Published as [DHW92].

Dongarra:1992:NCC

[DHW92] Jack J. Dongarra, Sven Hammarling, and James H. Wilkinson. Numerical considerations in computing invariant subspaces. SIAM Journal on Matrix Analysis and Applications, 13(1):145– 161, 1992. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Dubrulle:1971:IA

[DMW71] A. Dubrulle, R. S. Martin, and J. H. Wilkinson. The implicit QL algorithm. In Wilkinson and Reinsch [WR71], pages 241– 248. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Dongarra:1981:IAC

[DMW81] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. Technical Report ANL-81-43, Applied Mathematics Division, Argonne National Laboratory, Argonne, Illinois 60439, USA, 1981. Published as [DMW83].

Dongarra:1983:IAC

[DMW83] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. SIAM Journal on Numerical Analysis, 20(1):23–45, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davies:1959:ACE

[DW59] D. W. Davies and J. H. Wilkinson. The Automatic Computing Engine at the National Physical Laboratory. Nature, Lond., 183: 22–23, 1959.

Evans:1974:SNM

[Eva74] D. J. Evans, editor. Software for Numerical Mathematics. Academic Press, New York, NY, USA, 1974. ISBN 0-12-243750-0. LCCN QA297 .S591. Proceedings of the IMA Conference, Loughborough University of Technology, 1973.

Evans:1976:IJH

[Eva76] Christopher Evans. Interview with J. H. Wilkinson, page ?? Number 10 in Pioneers of Computing, 60-Minute Recordings of Interviews. Science Museum, London, UK, 1976.

Fox:1957:MCM

[FGM^{+}57] L. Fox, E. T. Goodwin, J. G. L. Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of Notes on applied science. HMSO, London, UK, 1957. vi + 128 pp. LCCN QA297 .T4.

Fox:1970:MCM

[FGM^{+}70] L. Fox, E. T. Goodwin, J. G. L.

Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of NPL Notes on Applied Science. HMSO, London, UK, second edition, 1970. ISBN 0-11-480021-9. vii + 169 pp. LCCN QA297 .T4 1970.

Fox:1948:NSA

[FHW48a] L. Fox, H. D. Huskey, and J. H.

Wilkinson. Notes on the solution of algebraic linear simultaneous equations. Quarterly Journal of Mechanics and Applied Mathematics, 1:149–173, 1948. CO-DEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1948:SAL

[FHW48b] L. Fox, H. D. Huskey, and J. H.

Wilkinson. The solution of algebraic linear simultaneous equations by punched card methods. Maths, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1948.

Fox:1987:JHW

[Fox87a] L. Fox. James Hardy Wilkinson, 27 September 1919–5 October 1986. Biographical Memoirs of Fellows of the Royal Society, 33: 669–708, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

URL http://www.jstor.org/
stable/769967;   https://
royalsocietypublishing.org/

                                        doi/epdf/10.1098/rsbm.1987.
                                        0024.

Fox:1987:ODJ

[Fox87b] Leslie Fox. Obituary: Dr. J. H. Wilkinson, F.R.S. The Computer Journal, 30(1):1, 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_30/Issue_
                                        01/tiff/1.tif.

Ferreira:2020:EWM

[FP20] Carla Ferreira and Beresford Parlett. Eigenpairs of Wilkinson matrices. SIAM Journal on Matrix Analysis and Applications, 41(3):1388–1415, ???? 2020. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Fox:1975:F

[FW75] L. Fox and J. H. Wilkinson. Foreward. In The NAG Library Manual, pages 1–3. The Numerical Algorithms Group Ltd, Wilkinson House, Jordan Hill Road, Oxford OX2 8DR, UK, 1975.

Golub:1984:SNA

[Gol84] Gene H. Golub, editor. Studies in Numerical Analysis, volume 24 of Studies in mathematics. Mathematical Association of America, Washington, DC, 1984. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). x + 415 pp. LCCN QA297 .S83 1984.

Grau:1966:BRB

[Gra66] A. A. Grau. Book review:

Rounding Errors in Algebraic Processes (J. H. Wilkinson). SIAM Review, 8(3):397–398, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golub:1970:LAS

[GUW70] G. H. Golub, R. Underwood, and J. H. Wilkinson. The Lanczos algorithm for the symmetric Ax = λBx problem. Technical report 142, Computer Science Department, Stanford University, Stanford, CA, USA, 1970.

Golub:1972:LAS

[GUW72] G. H. Golub, R. Underwood, and J. H. Wilkinson. The Lanczos algorithm for the symmetric Ax = λBx problem. Technical Report STAN-CS-72-270, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1972. 24 pp.

Golub:1966:IRL

[GW66a] G. H. Golub and J. H. Wilkinson. Iterative refinement of least square solution. In Proceedings of the IFIP Congress 65, Volume 2, pages 606–607. International Federation for Information Processing (IFIP), 1966.

Golub:1966:NIR

[GW66b] G. H. Golub and J. H. Wilkinson. Note on the iterative refinement of least squares solution.

Numerische Mathematik, 9(2): 139–148, December 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1975:ICE

[GW75] G. H. Golub and J. H. Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. Technical Report STAN-CS-75-478, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1975. Published as [GW76].

Golub:1976:ICE

[GW76] G. H. Golub and J. H. Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. SIAM Review, 18(4):578–619, 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammarling:2019:CCJ

[HH19a] Sven Hammarling and Nick Higham. Celebrating the centenary of James H. Wilkinson’s birth. Web site and conference, May 29–30, 2019. URL https:/

/nla-group.org/2019/01/21/
celebrating-the-centenary-
of-james-h-wilkinsons-birth/

.

Hammarling:2019:WQ

[HH19b] Sven Hammarling and Nick Higham. Wilkinson quotes. Web site, May 13, 2019.

URL https://nla-group.org/
2019/05/13/wilkinson-quotes/

.

Hammarling:2019:CJH

[HH19c] Sven J. Hammarling and Nicholas J.

Higham. Celebrating James Hardy Wilkinson. Mathematics Today, 55(4):131, 2019. ISSN 1361-2042.

Hammarling:2019:HAN

[HHT19] Sven Hammarling, Nick Higham, and Françoise Tisseur. Highlights of Advances in Numerical Linear Algebra Conference. Web site., June 19, 2019. URL https:/

/nla-group.org/2019/06/19/
highlights-of-advances-in-
numerical-linear-algebra-
conference/.

Hartree:1948:DCM

[HNW^{+}48] D. R. Hartree, M. H. A. Newman, M. V. Wilkes, F. C. Williams, J. H. Wilkinson, and A. D. Booth. A discussion on computing machines. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195 (1042):265–287 (2 plates), December 22, 1948. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

URL http://www.jstor.org/
stable/98328.

Householder:1970:NGI

[HVW70] Alston S. Householder, Richard S.

Varga, and James H. Wilkinson. A note on Gerschgorin’s inclusion theorem for eigenvalues of matrices. Numerische Mathematik, 16:141–144, 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hammarling:1976:PBL

[HW76a] S. Hammarling and J. H. Wilkinson. The practical behaviour of linear iterative methods with particular reference to S.O.R. Technical Report NAC 69, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1976.

Hartnell:1976:PFM

[HW76b] B. L. Hartnell and H. C. Williams, editors. Proceedings of the Fifth Manitoba Conference on Numerical Mathematics, October 1–4, 1975, number 16 in Congressus numerantium. Utilitas Mathematica Publishers, Winnipeg, Manitoba, Canada, 1976. ISBN 0-919628-16-8. LCCN QA297 .M33 1975.

Hammarling:1980:LSA

[HW80] S. Hammarling and J. H. Wilkinson. On linear systems arising from finite difference approximations to elliptic differential equations. Technical Report DNACS 34/80, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1980.

I:1970:BRJ

[I.70] E. I. Book review: J. A. Wilkinson, Rundungsfehler, translated from English into German by G. Goos, Springer-Verlag, New York, 1969, x + 208 pp., 21 cm. Price $3.70 (paperbound). Mathematics of Computation, 24(109):229–230, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic).  URL http://www.
jstor.com/stable/2004899.

ICM:1974:PIC

[ICM75] Proceedings of the International Congress of Mathematicians, Vancouver, Canada, August 21– 29, 1974. Canadian Mathematical Congress, Montréal, Québec, Canada, 1975. ISBN 0-8218-3059-7. LCCN QA3 .A572.

Iserles:1987:SAN

[IP87] A. Iserles and M. J. D. Powell, editors. State of the Art in Numerical Analysis. Proceedings of the Joint IMA/SIAM Conference held at the University of Birmingham, 14–18 April 1986, volume 9 of The Institute of Mathematics and Its Applications conference series; new series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-853614-3. LCCN QA297.J65 1986. URL http:

//www.gbv.de/dms/hbz/toc/
ht002967923.pdf;   http://
zbmath.org/?q=an:0611.00024.

Jacobs:1976:SAN

[Jac76] David A. H. Jacobs, editor.

The state of the art in numerical analysis: Proceedings of the Conference on the State of the Art in Numerical Analysis held at the University of York, April 12th–15th, 1976. Academic Press, New York, NY, USA, 1976. ISBN 0-12-378650-9. LCCN QA297 .C6461 1976.

Jacobs:1978:NSN

[Jac78] D. A. H. Jacobs, editor. Numerical software, needs and availability: proceedings of the Conference on Applications of Numerical Software, Needs and Availability, held at the University of Sussex, September 19th–22nd, 1977. Academic Press, New York, NY, USA, 1978. ISBN 0-12-378660-6. LCCN QA297 C74 1977. UK£11.60.

Jacobson:1962:SEA

[JRW62] A. W. Jacobson, S. Rosen, and J. H. Wilkinson. Scientific and engineering applications of digital computers. In H. D. Huskey and G. A. Korn, editors, Computer Handbook, pages 21:2– 21:13. McGraw-Hill, New York, 1962. Section 2.1.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Moler:2022:CCC

[Mol22] Cleve Moler. Cleve’s corner:

Christian Reinsch, Roland Bulirsch, and the SVD. Math-Works Web site., October 23,

2022. URL http://feeds.
feedburner.com/mathworks/
moler.

Martin:1965:SDPb

[MPW65] R. S. Martin, G. Peters, and J. H. Wilkinson. Symmetric decomposition of a positive definite matrix. Numerische Mathematik, 7(5):362–383, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 9–30].

Martin:1966:IRS

[MPW66] R. S. Martin, G. Peters, and J. H. Wilkinson. Iterative refinement of the solution of a positive definite system of equations. Numerische Mathematik, 8(3):203–216, May 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 31– 44].

Martin:1970:ARH

[MPW70] R. S. Martin, G. Peters, and J. H. Wilkinson. The QR algorithm for real Hessenberg matrices. Numerische Mathematik, 14(3):219–231, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 359– 371].

Martin:1971:ARH

[MPW71a] R. S. Martin, G. Peters, and J. H. Wilkinson. The QR algorithm for real Hessenberg matrices. In Wilkinson and Reinsch [WR71], pages 359– 371. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:IRS

[MPW71b] R. S. Martin, G. Peters, and J. H. Wilkinson. Iterative refinement of the solution of a positive definite system of equations. In Wilkinson and Reinsch [WR71], pages 31–44. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:SDPa

[MPW71c] R. S. Martin, G. Peters, and J. H. Wilkinson. Symmetric decomposition of a positive definite matrix. In Wilkinson and Reinsch [WR71], pages 9– 30. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1968:HTS

[MRW68] R. S. Martin, C. Reinsch, and J. H. Wilkinson. Householder’s tridiagonalization of a symmetric matrix. Numerische Mathematik, 11(3):181–195, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 212–226].

Martin:1970:ABS

[MRW70] R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR algorithm for band symmetric matrices. Numerische Mathematik, 16(2):85–92, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See also [WR71, pp. 266–272].

Martin:1971:ABS

[MRW71a] R. S. Martin, C. H. Reinsch, and J. H. Wilkinson. The QR algorithm for band symmetric matrices. In Wilkinson and Reinsch [WR71], pages 266– 272. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:HTS

[MRW71b] R. S. Martin, C. H. Reinsch, and J. H. Wilkinson. Householder’s tridiagonalization of a symmetric matrix. In Wilkinson and Reinsch [WR71], pages 212– 226. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Metropolis:1963:PFS

[MTTT63] N. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Experimental arithmetic, high speed computing and mathematics: Proceedings of the fifteenth Symposium in Applied Mathematics of the American Mathematical Society held in Chicago, Illinois, April 12–14, 1962 and Atlantic City, New Jersey, April 16-19, 1962, volume 15. American Mathematical Society, Providence, RI, USA, 1963. LCCN QA297 .S987 1962.

Martin:1965:SDPa

[MW65] R. S. Martin and J. H. Wilkinson. Symmetric decomposition of positive definite band matrices. Numerische Mathematik, 7 (5):355–361, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 50– 56].

Martin:1967:SSU

[MW67] R. S. Martin and J. H. Wilkinson. Solution of symmetric and unsymmetric band equations and the calculation of eigenvectors of band matrices. Numerische Mathematik, 9(4): 279–301, February 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 70– 92].

Martin:1968:RSE

[MW68a] R. S. Martin and J. H. Wilkinson. Reduction of the symmetric eigenproblem Ax = λBx and related problems to standard form. Numerische Mathematik, 11(2):99–110, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 303–314].

Martin:1968:SRG

[MW68b] R. S. Martin and J. H. Wilkinson. Similarity reduction of a general matrix to Hessenberg form. Numerische Mathematik, 12(5):349–368, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 339– 358].

Martin:1968:IA

[MW68c] Roger S. Martin and J. H.

Wilkinson. The implicit QL algorithm. Numerische Mathematik, 12(5):377–383, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 241–248], where A. A. Dubrulle and is listed as the lead author.

Martin:1968:MAC

[MW68d] Roger S. Martin and J. H.

Wilkinson. The modified LR algorithm for complex Hessenberg matrices. Numerische Mathematik, 12(5):369–376, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 396–403].

Martin:1971:MAC

[MW71a] R. S. Martin and J. H. Wilkinson. The modified LR algorithm for complex Hessenberg matrices. In Wilkinson and Reinsch [WR71], pages 396– 403. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:RSE

[MW71b] R. S. Martin and J. H. Wilkinson. Reduction of the symmetric eigenproblem Ax = λBx and related problems to standard form. In Wilkinson and Reinsch [WR71], pages 303– 314. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:SRG

[MW71c] R. S. Martin and J. H. Wilkinson. Similarity reduction of a general matrix to Hessenberg form. In Wilkinson and Reinsch [WR71], pages 339– 358. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:SSU

[MW71d] R. S. Martin and J. H. Wilkinson. Solution of symmetric and unsymmetric band equations and the calculation of eigenvectors of band matrices. In Wilkinson and Reinsch [WR71], pages 70–92. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Martin:1971:SDPb

[MW71e] R. S. Martin and J. H. Wilkinson. Symmetric decomposition of positive definite band matrices. In Wilkinson and Reinsch [WR71], pages 50–56. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Meek:1981:PTM

[MW81] D. S. Meek and Hugh C. Williams, editors. Proceedings of the tenth Manitoba Conference on Numerical Mathematics and Computing, October 1–4, 1980, volume 30–31 of Congressus Numerantium. Utilitas Mathematica Publishers, Winnipeg, Manitoba, Canada, 1981. ISBN 0-919628-30-3 (vol. 1) 0-919628-31-1 (vol. 2). ISSN 0384-9864.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Nash:1985:BC

[NW85] J. C. Nash and J. H. Wilkinson. The birth of a computer. BYTE Magazine, 10(2):177–??, ???? 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Oettli:1964:ASL

[OPW64] W. Oettli, W. Prager, and J. H.

Wilkinson. Admissible solutions of linear systems with not sharply defined coefficients. Research Paper RZ-166, IBM Research Division, IBM Zurich Research Laboratory, Rüschlikon, Zurich, Switzerland, 1964. Published as [OPW65].

Oettli:1965:ASL

[OPW65] W. Oettli, W. Prager, and J. H.

Wilkinson. Admissible solutions of linear systems with not sharply defined coefficients. SIAM J. Num. Anal., 2:291–299, 1965.

Olver:1982:PEB

[OW82] F. W. J. Olver and J. H. Wilkinson. A posteriori error bounds for Gaussian elimination. IMA Journal of Numerical Analysis, 2(4):377–406, 1982. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Parlett:1966:BRB

[Par66] B. N. Parlett. Book review: The Algebraic Eigenvalue Problem (J. H. Wilkinson). SIAM Review, 8(4):543–545, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Partlett:1972:BRB

[Par72] B. Partlett. Book review: Handbook for Automatic Computation, Vol. II, Linear Algebra (J. H. Wilkinson and C. Reinsch). SIAM Review, 14(4):658–661, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1987:CJH

[Par87] Beresford N. Parlett. A contribution of J. H. Wilkinson to numerical analysis. Technical Report PAM-372, Center for Pure and Applied Mathematics, University of California, Berkeley, Berkeley, CA, USA, April 1987.

Parlett:1990:CJH

[Par90] B. N. Parlett. The contribution of J. H. Wilkinson to numerical analysis. In Nash [Nas90], pages 17–30. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Powell:1977:CDJ

[Pow77] M. J. D. Powell, editor. The Contribution of Dr. J. H. Wilkinson to numerical analysis: Proceedings of a symposium … held in London on July 6th, 1977, number 19 in IMA Symposium Proceedings. Institute of Mathematics and its Applications, Southend-on-Sea, UK, 1977. LCCN QA297.C66 1978.

Peters:1969:EBS

[PW69] G. Peters and J. H. Wilkinson. Eigenvalues of Ax = λBx with band symmetric A and B. The Computer Journal, 12(4):398–404, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120397.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/398.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/399.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/400.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/401.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/402.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/403.tif;   http:/
/www3.oup.co.uk/computer_

                                        journal/hdb/Volume_12/Issue_
                                        04/tiff/404.tif.

Peters:1970:GE

[PW70a] G. Peters and J. H. Wilkinson. Ax = λBx and the generalized eigenproblem. SIAM Journal on Numerical Analysis, 7:479– 492, 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1970:ERC

[PW70b] G. Peters and J. H. Wilkinson. Eigenvectors of real and complex matrices by LR and QR triangularization. Numerische Mathematik, 16(3):181–204, December 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 372–395].

Peters:1970:LSP

[PW70c] G. Peters and J. H. Wilkinson. The least squares problem and pseudo-inverses. The Computer Journal, 13(3):309–316, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/130309.sgm.abs.html;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/309.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/310.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_13/Issue_
                                        03/tiff/311.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/312.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/313.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/314.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/315.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/316.tif.

Peters:1971:CSE

[PW71a] G. Peters and J. H. Wilkinson. The calculation of specified eigenvectors by inverse iteration. In Wilkinson and Reinsch [WR71], pages 418– 439. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Peters:1971:ERC

[PW71b] G. Peters and J. H. Wilkinson.

Eigenvectors of real and complex matrices by LR and QR triangularizations. In Wilkinson and Reinsch [WR71], pages 372– 395. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Peters:1971:PPA

[PW71c] G. Peters and J. H. Wilkinson. Practical problems arising in the solution of polynomial equations. Journal of the Institute of Mathematics and its Applications, 8:16–35, 1971. CO-DEN JMTAA8. ISSN 0020-

Peters:1974:ACE

[PW74a] G. Peters and J. H. Wilkinson. Accuracy of computed eigensystems and invariant subspaces. In Scaife [Sca74], pages 115–135. ISBN 0-12-621150-7. LCCN QA297 .S86. URL http:

                                        //catalog.hathitrust.org/
                                        Record/000575557. Festschrift

in honour of C. Lanczos commissioned by the Royal Irish Academy.

Peters:1974:APD

[PW74b] G. Peters and J. H. Wilkinson. On an algorithm for polynomial deflation by Broyden and Ford. Technical Report NAC 55, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1974.

Peters:1974:SAS

[PW74c] G. Peters and J. H. Wilkinson. Some algorithms for the solution of the generalized symmetric eigenvalue problem Au = λBu. Unpublished, Circa 1974. National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK., 1974.

Peters:1975:SGJ

[PW75] G. Peters and J. H. Wilkinson. On the stability of Gauss–Jordan elimination with pivoting. Communications of the ACM, 18(1):20–24, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Peters:1979:III

[PW79] G. Peters and J. H. Wilkinson. Inverse iteration, illconditioned equations and Newton’s method. SIAM Review, 21(3):339–360, 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1965:EDCa

[Ral65a] L. B. Rall, editor. Error in Digital Computation, volume 1. Wiley, New York, NY, USA, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Rall:1965:EDCb

[Ral65b] L. B. Rall, editor. Error in Digital Computation, volume 2. Wiley, New York, NY, USA, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Reinsch:2023:SVD

[RR23] Christian Reinsch and Mathias Richter. Singular value decomposition in extended double precision arithmetic. Numerical Algorithms, 93(3):1137–1155, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/
10.1007/s11075-022-01459-9.

Rollett:1961:ESC

[RW61] J. S. Rollett and J. H. Wilkinson. An efficient scheme for the co-diagonalization of a symmetric matrix by Givens’ method in a computer with a twolevel store. The Computer Journal, 4(2):177–180, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/4/2/177.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_02/040177.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/177.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/178.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/179.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
02/tiff/180.tif.

Rollett:1962:LEG

[RW62] J. S. Rollett and J. H. Wilkinson. Letter to the editor: [Givens and Householder processes]. The Computer Journal, 4(4):279, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
cgi/reprint/4/4/279;   http:
//comjnl.oxfordjournals.org/
content/4/4/279.full.pdf+
html.

Scaife:1974:SNA

[Sca74] B. K. P. (Brendan Kevin Patrick)

Scaife, editor. Studies in numerical analysis: papers in honour of Cornelius Lanczos. Academic Press, New York, NY, USA, 1974. ISBN 0-12-621150-7. xxii + 333 pp. LCCN QA297 .S86. URL http:

//catalog.hathitrust.org/
Record/000575557. Published

for the Royal Irish Academy.

Symm:1980:REBa

[SW80a] H. J. Symm and J. H. Wilkinson. Realistic error bounds for a simple eigenvalue and its associated eigenvector. Technical Report STAN-CS-80-787, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1980. Published as [SW80b].

Symm:1980:REBb

[SW80b] H. J. Symm and J. H. Wilkinson.

Realistic error bounds for a simple eigenvalue and its associated eigenvector. Numerische Mathematik, 35(2):113–126, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Symm:1981:EBC

[SW81a] H. J. Symm and J. H. Wilkinson. Error bounds for computed invariant subspaces. Research Report 81-02, Seminar für Angewandte Mathematik, Eidgenössische Technische Hochschule, ETH-Zentrum CH-8092 Zurich, Switzerland, 1981. Lecture presented at a Symposium in honour of H. Rutishauser, October 15–17, 1980.

Symm:1981:ADM

[SW81b] Hilary J. Symm and J. H. Wilkinson. Almost defective matrices. In Meek and Williams [MW81], pages 89–131. ISBN 0-919628-30-3 (vol. 1) 0-919628-31-1 (vol. 2). ISSN 0384-9864.

Turing:2005:TWL

[TW05] Alan M. Turing and James H. Wilkinson. The Turing– Wilkinson lecture series (1946– 1947). In Copeland [Cop05], page ?? ISBN 0-19-856593-3 (hardcover). LCCN QA75 .A43 2005. URL http://

ukcatalogue.oup.com/product/
9780198565932.do;   http:/
/www.oxfordscholarship.com/
oso/public/content/maths/
9780198565932/toc.html. Edited

with an introduction by B. Jack Copeland.

Turing:2012:TWL

[TW12] Alan M. Turing and James H. Wilkinson. The Turing– Wilkinson lecture series (1946– 7) (edited with an introduction by B. Jack Copeland). In Copeland [Cop12], pages 459– 527. ISBN 0-19-960915-2 (paperback). LCCN ???? UK £14.99.

Watkins:2011:FA

[Wat11] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118(5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/info/
10.4169/amer.math.monthly.
118.05.387.

Williams:1989:EBC

[WCK89] M. R. Williams and Martin Campbell-Kelly, editors. The Early British Computer Conferences, volume 14 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-23136-0. xvi + 508 pp. LCCN QA76.17 .E171 1989.

Wilkinson:1948:ACE

[Wil48a] J. H. Wilkinson. The Automatic Computing Engine at the National Physical Laboratory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195(1042):285–286, December 22, 1948. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

URL http://www.jstor.org/
stable/98328.

Wilkinson:1948:PRA

[Wil48b] J. H. Wilkinson. Progress report on the Automatic Computing Engine. Report MA/17/1024, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, April 1948. ???? pp.

URL http://www.alanturing.
net/turing_archive/archive/
l/l10/L10-001.html.

Wilkinson:1951:PGR

[Wil51a] J. H. Wilkinson. Programmes for the GPO Radar Data Analyser. Maths, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1951.

Wilkinson:1951:PCP

[Wil51b] J. H. Wilkinson. Programming and coding for the pilot model of the ACE. Ma report, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1951.

Wilkinson:1951:RPM

[Wil51c] J. H. Wilkinson. Report on the pilot model of the Automatic Computing Engine. II. the logical design of the pilot model. Maths division and electronics section, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1951.

Wilkinson:1952:PCP

[Wil52] J. H. Wilkinson. Programming and coding for the pilot model of the ACE. Ma Report No. 22, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1952. Reissue.

Wilkinson:1954:CLR

[Wil54a] J. H. Wilkinson. The calculation of the latent roots and vectors of matrices on the Pilot Model of the ACE. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 50(4):536–566, October 1954. CODEN PCPSA4. ISSN 0008-1981.

Wilkinson:1954:LAP

[Wil54b] J. H. Wilkinson. Linear algebra on the Pilot ACE. In Automatic Digital Computation, pages 129– 136. HMSO, London, UK, 1954. Proceedings of Symposium at the National Physical Laboratory, 1953.

Wilkinson:1954:PA

[Wil54c] J. H. Wilkinson. The Pilot ACE. In Automatic Digital Computation, pages 5–14. HMSO, London, UK, 1954. Proceedings of Symposium at the National Physical Laboratory, 1953. Reprinted in [BN71, pp. 193–199] and in [WCK89, pp. 219–228].

Wilkinson:1955:ASO

[Wil55a] J. H. Wilkinson. An assessment of the system of optimum coding used on the Pilot ACE at the National Physical Laboratory. Philosophical transactions of the Royal Society of London Series A, 248: 253–281, 1955. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Wilkinson:1955:ECM

[Wil55b] J. H. Wilkinson. Electronic computing machines and their uses. J. scient. Instrum., 32:409–415, 1955.

Wilkinson:1955:RVU

[Wil55c] J. H. Wilkinson. Report on visit to U.S.A., 17 September– 14 December 1954. Maths, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1955.

Wilkinson:1955:UIM

[Wil55d] J. H. Wilkinson. The use of iterative methods for finding the latent roots and vectors of matrices. Mathematical Tables and Other Aids to Computation, 9(52):184–191, October 1955. CODEN MTTCAS. ISSN 0891-6837.

Wilkinson:1957:BPD

[Wil57a] J. H. Wilkinson. British progress in digital computer design. Automn Prog., 2:5–9, 36, 1957.

Wilkinson:1957:CEC

[Wil57b] J. H. Wilkinson. The calculation of the eigenvectors of codiagonal matrices produced by the Givens and Lanczos processes. In Conference Proceedings, pages 112:1–112:15. Weapons Research Establishment, Salisbury, Australia, 1957.

Wilkinson:1957:CFG

[Wil57c] J. H. Wilkinson. Computing facilities in Great Britain. J. Inst. Petrol., 43:101–107, 1957.

Wilkinson:1957:EUD

[Wil57d] J. H. Wilkinson. Example of the use of DEUCE. In Modern Computing Methods, 1st Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 116–122. HMSO, London, UK, 1957.

Wilkinson:1957:HPD

[Wil57e] J. H. Wilkinson. Hyperbolic partial differential equations. In Modern Computing Methods, 1st Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 61–71. HMSO, London, UK, 1957.

Wilkinson:1957:LRM

[Wil57f] J. H. Wilkinson. Latent roots of matrices. In Modern Computing Methods, 1st Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 27–35. HMSO, London, UK, 1957.

Wilkinson:1957:LEM

[Wil57g] J. H. Wilkinson. Linear equations and matrices. In Modern Computing Methods, 1st Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 12–19. HMSO, London, UK, 1957.

Wilkinson:1957:LOW

[Wil57h] J. H. Wilkinson. Loading on an oscillating wing of infinite aspect ratio in a subsonic stream. ARC, 19:344, 1957. Comp. 125.

Wilkinson:1958:CEM

[Wil58a] J. H. Wilkinson. The calculation of eigenvectors by the method of Lanczos. The Computer Journal, 1(3):148–152, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_

journal/hdb/Volume_01/Issue_
03/010148.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/148.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/149.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/150.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/151.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/152.tif.

Wilkinson:1958:CEC

[Wil58b] J. H. Wilkinson. The calculation of the eigenvectors of codiagonal matrices. The Computer Journal, 1(2):90–96, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/010090.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/90.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/
91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/92.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/93.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
02/tiff/94.tif;   http://
www3.oup.co.uk/computer_journal/

hdb/Volume_01/Issue_02/tiff/
95.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/96.
tif.

Wilkinson:1958:MPN

[Wil58c] J. H. Wilkinson. Matrix programs at the National Physical Laboratory. J. ACM, 5:113–114, 1958.

Wilkinson:1958:SRM

[Wil58d] J. H. Wilkinson. Stability of the reduction of a matrix to almost triangular and triangular forms by elementary similarity transformations. Mathematics Division No. 22, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1958. Reprinted 1962. Published as [Wil59c].

Wilkinson:1959:EZIa

[Wil59a] J. H. Wilkinson. The evaluation of the zeros of ill-conditioned polynomials. I. Numerische Mathematik, 1:150–166, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1959:EZIb

[Wil59b] J. H. Wilkinson. The evaluation of the zeros of ill-conditioned polynomials. II. Numerische Mathematik, 1:167–180, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1959:SRM

[Wil59c] J. H. Wilkinson. Stability of the reduction of a matrix to almost triangular and triangular forms by elementary similarity transformations. J. ACM, 6(3):336– 359, 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1960:EAF

[Wil60a] J. H. Wilkinson. Error analysis of floating-point computation. Numerische Mathematik, 2(1):319–340, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/
10.1007/BF01386233.

Wilkinson:1960:HMS

[Wil60b] J. H. Wilkinson. Householder’s method for the solution of the algebraic eigenproblem. The Computer Journal, 3(1):23–27, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/030023.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/23.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/
24.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/25.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/26.tif;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
01/tiff/27.tif.

Wilkinson:1960:MSL

[Wil60c] J. H. Wilkinson. On methods for solving linear equations. In Proceedings of UNESCO Conference on Information Processing, 1959, pages 108–109. Butterworth, London, UK, 1960.

Wilkinson:1960:REA

[Wil60d] J. H. Wilkinson. Rounding errors in algebraic processes. In Proceedings of UNESCO Conference on Information Processing, 1959, pages 44–53. Butterworth, London, UK, 1960.

Wilkinson:1961:EA

[Wil61a] J. H. Wilkinson. Error analysis.

In Modern Computing Methods, 2nd Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 41–

52. HMSO, London, UK, 1961.

Wilkinson:1961:EAD

[Wil61b] J. H. Wilkinson. Error analysis of direct methods of matrix inversion. Journal of the ACM, 8 (3):281–330, 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1961:HPD

[Wil61c] J. H. Wilkinson. Hyperbolic partial differential equations. In Modern Computing Methods, 2nd Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 101–

111. HMSO, London, UK, 1961.

Wilkinson:1961:LEM

[Wil61d] J. H. Wilkinson. Linear equations and matrices. In Modern Computing Methods, 2nd Edition, National Physical Laboratory, Notes on Applied Science, Volume 16, pages 13–33. HMSO, London, UK, 1961.

Wilkinson:1961:RA

[Wil61e] J. H. Wilkinson. Remarks on Algorithms 2 and 3. Communications of the ACM, 4(3): 153, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkinson:1961:REB

[Wil61f] J. H. Wilkinson. Rigorous error bounds for computed eigensystems. The Computer Journal, 4(3):230–241, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/040230.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/230.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/231.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/232.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/233.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/234.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/235.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/236.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/237.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/238.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/239.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/240.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
03/tiff/241.tif.

Wilkinson:1962:BRB

[Wil62a] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, by E. Durand (Paris: Masson et Cie., 1960, pp. 328, 65 NF). The Computer Journal, 5(1):32, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/5/1/28.full.pdf;
http://comjnl.oxfordjournals.
org/content/5/1/32.full.pdf+
html.

Wilkinson:1962:CESa

[Wil62b] J. H. Wilkinson. Calculation of the eigenvalues of a symmetric tridiagonal matrix by the method of bisection. Numerische Mathematik, 4:362– 367, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:CESb

[Wil62c] J. H. Wilkinson. Calculation of the eigenvectors of a symmetric tridiagonal matrix by inverse iteration. Numerische Mathematik, 4:368–376, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:EAE

[Wil62d] J. H. Wilkinson. Error analysis of eigenvalue techniques based on orthogonal transformations. Journal of the Society for Industrial and Applied Mathematics, 10:162–195, 1962. CODEN JSI-MAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Wilkinson:1962:HMS

[Wil62e] J. H. Wilkinson. Householder’s method for symmetric matrices. Numerische Mathematik, 4:354– 361, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:IEM

[Wil62f]  J. H. Wilkinson.  Instabil-

ity of the elimination method of reducing a matrix to tridiagonal form. The Computer Journal, 5(1):61–70, April 1962. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/050061.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/61.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/63.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/65.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/67.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
01/tiff/69.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/
70.tif.

Wilkinson:1962:NQC

[Wil62g] J. H. Wilkinson. Note on the quadratic convergence of the cyclic Jacobi process. Numerische Mathematik, 4:296– 300, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:U

[Wil62h] James Hardy Wilkinson. [unknown]. ScD, Cambridge University, Cambridge, UK, 1962.

Wilkinson:1963:BRB

[Wil63a] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, Volume II, by E. Durand (Paris: Masson et Cie., 1961, pp. 445, 90 NF). The Computer Journal, 5(4):263, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/5/4/258.full.pdf;
                                         http://comjnl.oxfordjournals.
                                        org/content/5/4/263.full.
                                        pdf+html.

Wilkinson:1963:ELS

[Wil63b] J. H. Wilkinson. Errors in largescale numerical problems. Comput. Bull., 6:124–125, 1963.

Wilkinson:1963:PRF

[Wil63c] J. H. Wilkinson. Plane rotations in floating-point arithmetic. In Metropolis et al. [MTTT63], pages 185–198. LCCN QA297 .S987 1962.

Wilkinson:1963:REA

[Wil63d] J. H. Wilkinson. Rounding Errors in Algebraic Processes, volume 32 of Notes on Applied Science. HMSO, London, UK, 1963. ISBN 0-486-67999-3 (Dover). vi + 161 pp. LCCN QA76.5 .W53 1964. Also published by Prentice-Hall, Englewood Cliffs, NJ, USA, 1964, translated into Polish as Bledy Zaokragleń w Procesach Algebraicznych by PWW, Warsaw, Poland, 1967 and translated into German as Rundungsfehler by Springer-Verlag, Berlin, Germany, 1969. Reprinted by Dover Publications, New York, 1994.

Wilkinson:1965:ARS

[Wil65a] J. H. Wilkinson. The QR algorithm for real symmetric matrices with multiple eigenvalues. The Computer Journal, 8(1):85–87, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080085.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/85.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
86.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/87.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/88.tif.

Wilkinson:1965:AEP

[Wil65b] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1965. ISBN 0-19-853403-5. xviii + 662 pp. LCCN QA218 .W686 1965.

Wilkinson:1965:CLP

[Wil65c] J. H. Wilkinson. The calculation of Lamé polynomials. The Computer Journal, 8(3):273–277, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/080273.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/273.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/274.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/275.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/276.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
03/tiff/277.tif.

Wilkinson:1965:CRA

[Wil65d]  J. H. Wilkinson.  Conver-

gence of the LR, QR and related algorithms. The Computer Journal, 8(1):77–84, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/080077.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/77.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
78.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/79.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_
01/tiff/81.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/83.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/84.tif.

Wilkinson:1965:EAT

[Wil65e] J. H. Wilkinson. Error analysis for transformations based on the use of matrices of the form I − 2ww^{H}. In Rall [Ral65b], pages 77–101. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Wilkes:1966:BRB

[Wil66a]  M. V. Wilkes.  Book re-

view: The Algebraic Eigenvalue Problem. The Computer Journal, 8(4):382, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/8/4/382.full.pdf+
html.

Wilkinson:1966:CEM

[Wil66b]  J. H. Wilkinson.  Calcula-

tion of eigensystems of matrices. In J. Walsh, editor, Numerical Analysis: An Introduction, pages 27–61. Academic Press, New York, NY, USA, 1966.

Wilkinson:1967:ADM

[Wil67a] J. H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Technical Report CS 59, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1967. Published as [Wil68a].

Wilkinson:1967:BZW

[Wil67b]  J. H. Wilkinson.  Bledy

Zaokragleń w Procesach Algebraicznych. (Polish) [Rounding errors in algebraic Processes]. PWW, Warszawa, Poland, 1967. ???? pp. Polish translation of [Wil63d].

Wilkinson:1967:BRB

[Wil67c] J. H. Wilkinson. Book review: The Matrix Analysis of Vibration, by R. E. D. Bishop, G. M. L. Gladwell, and S. Michaelson, 1965; 404 pages (London: Cambridge University Press, 100s). The Computer Journal, 10(1):77, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/10/1/74.full.pdf;
                                         http://comjnl.oxfordjournals.
                                        org/content/10/1/77.full.
                                        pdf+html.

Wilkinson:1967:SIC

[Wil67d] J. H. Wilkinson. The solution of ill-conditioned linear equations. In A. Ralston and H. Wilf, editors, Mathematical Methods for Digital Computers, Volume 2, pages 65–93. Wiley, New York, NY, USA, 1967.

Wilkinson:1967:SLA

[Wil67e] J. H. Wilkinson. Solution of linear algebraic equations and matrix problems by direct methods. In M. Klarer and G. A. Korn, editors, Digital Computer Users’ Handbook, pages 2:18– 2:55. McGraw-Hill, New York, 1967.

Wilkinson:1967:TAB

[Wil67f] J. H. Wilkinson. Two algorithms based on successive linear interpolation. Technical Report CS 60, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1967.

Wilkinson:1968:ADM

[Wil68a] J. H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Linear Algebra and its Applications, 1:1– 12, 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wilkinson:1968:BFE

[Wil68b] J. H. Wilkinson. The basic forms of error analysis of matrix processes. In Programmation en Mathématiques Numériques, pages 127–133. Éditions Centre Nat. Recherche Sci., Paris, France, 1968. Colloq. Int. Cent. Nat. Rech. Sci. 165, Besançon, Paris.

Wilkinson:1968:GCT

[Wil68c] J. H. Wilkinson. Global convergence of tridiagonal QR algorithm with origin shifts. Linear Algebra and its Applications, 1(3):409–420, July 1968. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379568900177.

Wilkinson:1968:PEA

[Wil68d]  J. H. Wilkinson.  A pri-

ori error analysis of algebraic processes. In I. G. Petrovsky, editor, Proceedings of International Congress of Mathematicians, pages 629–640. Izdatel’stvo Mir, Moscow, USSR, 1968.

Wilkinson:1968:SEA

[Wil68e] James Hardy Wilkinson. A survey of error analysis of matrix algorithms. Aplikace Matematiky, 13:93–102, 1968. CODEN APM-TAK. ISSN 0373-6725.

Wilkinson:1969:R

[Wil69a] J. H. Wilkinson. Rundungsfehler, volume 44 of Heidelberger Taschenbücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. x + 208 pp. German translation by G. Goos.

Wilkinson:1969:GCA

[Wil69b] James H. Wilkinson. Global convergence of QR algorithm (with discussion). In Proceedings of the IFIP Congress 1968, Edinburgh, Volume 1, pages 130–133. North-Holland, Amsterdam, The Netherlands, 1969.

Wilkinson:1970:APS

[Wil70a] J. H. Wilkinson. Algebraicheskaya problema sobstvennyh znachenii’. (Russian) [The Algebraic Eigenvalue Problem]. Nauka, Moscow, Russia, 1970. 564 pp. Russian translation of [Wil65b].

Wilkinson:1970:BRB

[Wil70b] J. H. Wilkinson. Book review: A Collection of Matrices for Testing Computational Algorithms, by Robert T. Gregory, D. L. Karney, 1969; 154 pages. (John Wiley and Sons Ltd., £4.75). The Computer Journal, 13(4):391, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/13/4/387.full.pdf;
 http://comjnl.oxfordjournals.
org/content/13/4/391.full.
pdf+html.

Wilkinson:1970:BRH

[Wil70c] J. H. Wilkinson. Book review: A Handbook of Numerical Matrix Inversion and Solution of Linear Equations, by Joan Westlake: pp. 171; 90s. (John Wiley and Sons, New York, 1968). Bulletin of the London Mathematical Society, 2(3):378–379, November 1970. CODEN LMSBBT. ISSN 0024-6093 (print), 1469-2120 (electronic). URL http:

                                        //blms.oxfordjournals.org/
                                        content/2/3/378.full.pdf.

Wilkinson:1970:EPW

[Wil70d] J. H. Wilkinson. Elementary proof of the Wielandt–Hoffman theorem and of its generalization. Technical Report CS 150, Computer Science Department, Stanford University, Stanford, California 94305, USA, 1970.

Wilkinson:1971:DNA

[Wil71a] J. H. Wilkinson. A discussion on numerical analysis of partial differential equations. Introductory remarks. Proc. R. Soc. Lond., A 323:153, 1971.

Wilkinson:1971:IPL

[Wil71b] J. H. Wilkinson. Introduction to Part I: Linear systems, least squares and linear programming. In Wilkinson and Reinsch [WR71], pages 1–8. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Wilkinson:1971:IPI

[Wil71c] J. H. Wilkinson. Introduction to Part II: The algebraic eigenvalue problem. In Wilkinson and Reinsch [WR71], pages 191– 201. ISBN 0-387-05414-6, 3-540-05414-6. LCCN QA251 .W67.

Wilkinson:1971:IR

[Wil71d] J. H. Wilkinson. Introductory remarks. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 323(??):153, June 8, 1971. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Wilkinson:1971:MEA

[Wil71e] J. H. Wilkinson. Modern error analysis. SIAM Review, 13 (4):548–568, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). The 1970 von Neumann lecture.

Wilkinson:1971:SCN

[Wil71f] J. H. Wilkinson. Some comments from a numerical analyst. Journal of the ACM, 18(2):137– 147, April 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1972:IIT

[Wil72a] J. H. Wilkinson. Inverse iteration in theory and in practice. In Symposia Matematico X, pages 361–379. Istituto Nazionale di Alta Matematica, Bologna, Italy, 1972. Monograph, Volume 10.

Wilkinson:1972:NMV

[Wil72b] J. H. Wilkinson. Note on matrices with a very illconditioned eigenproblem. Numerische Mathematik, 19(2): 176–178, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1972:NAN

[Wil72c] J. H. Wilkinson. Numerical analysis at the National Physical Laboratory. Bulletin of the Institute of Mathematics and its Applications, 8(9–10):297– 298, 1972. CODEN IMTABW. ISSN 0950-5628.

Wilkinson:1974:CEA

[Wil74a] J. H. Wilkinson. The classical error analysis for the solution of linear systems. Bulletin of the Institute of Mathematics and its Applications, 10 (5–6):175–180, 1974. CODEN IMTABW. ISSN 0950-5628.

Wilkinson:1974:HPE

[Wil74b] J. H. Wilkinson. Handbook p139 expanded or expounded. Unpublished, Circa 1974. National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK., 1974.

Wilkinson:1974:LAA

[Wil74c] J. H. Wilkinson. Linear algebra algorithms. In Evans [Eva74], pages 17–25 (discussion 27–28). ISBN 0-12-243750-0. LCCN QA297 .S591. With discussion by J. D. Beasley, E. L. Albasiny, O. Tingleff, Linda Hayes, M. G. Cox, M. J. D. Powell and J. H. Wilkinson.

Wilkinson:1974:NII

[Wil74d] J. H. Wilkinson. Note on inverse iteration and ill-conditioned eigensystems. Acta Univ. Carol. Math. Phys., 15(1–2):173–177, 1974.

Wilkinson:1974:NLA

[Wil74e] J. H. Wilkinson. Numerical linear algebra on digital computers. Bulletin of the Institute of Mathematics and its Applications, 10 (9–10):354–356, 1974. CODEN IMTABW. ISSN 0950-5628.

Wilkinson:1974:SVD

[Wil74f] J. H. Wilkinson. The singular value decomposition and minfit. Unpublished, Circa 1974. National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK., 1974.

Wilkinson:1974:TMA

[Wil74g] J. H. Wilkinson. Test matrices for algorithms for solving linear systems and inverting matrices. Unpublished, Circa 1974. National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK., 1974.

Wilkinson:1975:AHA

[Wil75a] J. H. Wilkinson. Alston S.

Householder award. Linear Algebra and its Applications, 11(1):1, ???? 1975. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379575901111.

Wilkinson:1975:IS

[Wil75b] J. H. Wilkinson. Invariant subspaces. In ICM’74 [ICM75], pages 443–448. ISBN 0-8218-3059-7. LCCN QA3 .A572.

Wilkinson:1975:PAN

[Wil75c] J. H. Wilkinson. The Pilot ACE at the NPL. The Radio and Electronic Engineer, 45(7):336–340, July 1975. CODEN RDEEA4. ISSN 0033-7722.

Wilkinson:1976:LDE

[Wil76] J. H. Wilkinson. Linear differential equations and Ax = λBx. In Hartnell and Williams [HW76b], pages 137–158. ISBN 0-919628-16-8. LCCN QA297 .M33 1975.

Wilkinson:1977:DSG

[Wil77a] J. H. Wilkinson. The differential system Bẋ = Ax and the generalized eigenvalue problem Au = λBu. Technical Report NAC 73, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1977.

Wilkinson:1977:SRA

[Wil77b] J. H. Wilkinson. Some recent advances in numerical linear algebra. In Jacobs [Jac76], pages 3– 23. ISBN 0-12-378650-9. LCCN QA297 .C6461 1976.

Wilkinson:1977:USP

[Wil77c] J. H. Wilkinson. The use of the single-precision residual in the solution of linear systems. Unpublished, Circa 1977. National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK., 1977.

Wilkinson:1978:AHA

[Wil78a] J. H. Wilkinson. The Alston S. Householder Award.

Numerische Mathematik, 29(4): 463, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1978:KCF

[Wil78b] J. H. Wilkinson. Kronecker’s canonical form and the QZ algorithm. Technical Report DNACS 10/78, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1978. Published as [Wil79].

Wilkinson:1978:LDE

[Wil78c] J. H. Wilkinson. Linear differential equations and Kronecker’s canonical form. In de Boor and Golub [dBG78], pages 231– 265. ISBN 0-12-208360-1. LCCN QA297 S994 1978.

Wilkinson:1978:NPS

[Wil78d] J. H. Wilkinson. Note on the practical significance of the Drazin inverse. Technical Report DNACS 13/79, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1978. Also produced as Stanford University, Computer Science Department Report STAN-CS-79-736. Published as [Wil82].

Wilkinson:1978:PNL

[Wil78e] J. H. Wilkinson. A problem in numerical linear algebra. In Powell [Pow77], pages 76–91. LCCN QA297.C66 1978.

Wilkinson:1978:SVD

[Wil78f] J. H. Wilkinson. Singular-value decomposition – basic aspects.

In Jacobs [Jac78], pages 109– 135. ISBN 0-12-378660-6. LCCN QA297 C74 1977. UK£11.60.

Wilkinson:1979:KCF

[Wil79] J. H. Wilkinson. Kronecker’s canonical form and the QZ algorithm. Linear Algebra and its Applications, 28:285–303, 1979. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Dedicated to the 75th birthday of A. S. Householder.

Wilkinson:1980:SRA

[Wil80a] J. H. Wilkinson. Some recent advances in numerical linear algebra. Yingyong Shuxue yu Jisuan Shuxue, 6:1–19, 1980. Translated from the English by Ji Guang Sun.

Wilkinson:1980:TWN

[Wil80b] J. H. Wilkinson. Turing’s work at the National Physical Laboratory and the construction of Pilot ACE, DEUCE, and ACE. In Metropolis et al. [MHR80], pages 101–114. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Wilkinson:1982:NPS

[Wil82] J. H. Wilkinson. Note on the practical significance of the Drazin inverse. In S. L. Campbell, editor, Recent Applications of Generalized Inverses, volume 66 of Res. Notes in Math., chapter 3, pages 82–99. Pitman Publishing, London, UK, 1982.

Wilkinson:1984:NMQ

[Wil84a] J. H. Wilkinson. On neighbouring matrices with quadratic elementary divisors. Numerische Mathematik, 44(1):1–21, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). In honour of the 60th birthday of F. L. Bauer.

Wilkinson:1984:SE

[Wil84b] J. H. Wilkinson. Sensitivity of eigenvalues. Util. Math., 25:5– 76, 1984. CODEN UTMADA. ISSN 0315-3681.

Wilkinson:1984:SAE

[Wil84c] J. H. Wilkinson. The state of the art in error analysis. Technical Memorandum 42, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, Illinois 60439, USA, 1984. Informal Proceedings of the Symposium on Computational Mathematics – State of the Art held at Argonne National Laboratory, September 20–21, 1984 in honour of the 65th birthday of James H. Wilkinson.

Wilkinson:1984:PP

[Wil84d] James H. Wilkinson. The perfidious polynomial. In Golub [Gol84], chapter 1, pages 1–28. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). LCCN QA297 .S83 1984. Awarded the Chauvenet Prize of the Mathematical Association of America.

Wilkinson:1985:SAE

[Wil85] J. H. Wilkinson. The state of the art in error analysis. NAG Newsletter, 2/85:5–28, 1985. Invited lecture for the NAG 1984 Annual General Meeting.

Wilkinson:1986:EAR

[Wil86a] J. H. Wilkinson. Error analysis revisited. Bulletin of the Institute of Mathematics and its Applications, 22(11–12):192– 200, 1986. CODEN IMTABW. ISSN 0950-5628. Invited lecture at Lancaster University in honour of C. W. Clenshaw, 1985.

Wilkinson:1986:ICN

[Wil86b] J. H. Wilkinson. Ill-condition in numerical linear algebra. Congressus Numerantium, 51:59–81, 1986. ISSN 0384-9864. Proceedings of the fifteenth Manitoba conference on numerical mathematics and computing (Winnipeg, Manitoba, 1985).

Wilkinson:1986:SEI

[Wil86c] J. H. Wilkinson. Sensitivity of eigenvalues II. Util. Math., 30: 243–286, 1986. CODEN UT-MADA. ISSN 0315-3681.

Wilkinson:1987:EP

[Wil87a]  J. H. Wilkinson.  Eigen-

value problems. In Iserles and Powell [IP87], pages 1–39. ISBN 0-19-853614-3. LCCN QA297.J65 1986. URL http:

//www.gbv.de/dms/hbz/toc/
ht002967923.pdf;   http://
zbmath.org/?q=an:0611.00024.

Proceedings of the IMA/SIAM Conference, University of Birmingham, 1986.

Wilkinson:1987:TF

[Wil87b] J. H. Wilkinson. On a theorem of Feingold. Linear Algebra and its Applications, 88/89:13– 30, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Issue in honour of the 65th birthday of J. H. Wilkinson.

Wilkinson:1987:SCN

[Wil87c] J. H. Wilkinson. Some comments from a numerical analyst. In Ashenhurst [Ash87], page ?? ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987.

Wilkinson:1988:AEP

[Wil88] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Monographs on Numerical Analysis. The Clarendon Press, Oxford University Press, New York, 1988. ISBN 0-19-853418-3. xviii + 662 pp. Oxford Science Publications.

Wilkinson:1994:REA

[Wil94] J. H. Wilkinson. Rounding errors in algebraic processes. Dover, New York, NY, USA, 1994. ISBN 0-486-67999-3. viii + 161 pp. Reprint of the 1963 original [Prentice-Hall, Englewood Cliffs, NJ; MR0161456 (28 #4661)].

Wilkinson:2005:PAN

[Wil05] James H. Wilkinson. The Pilot ACE at the National Physical Laboratory. In Copeland [Cop05], pages 93–105. ISBN 0-19-856593-3 (hardcover). LCCN QA75 .A43 2005. URL http:

                                        //www.ingentaconnect.com/
                                        content/oso/2353639/2005/
                                        00000001/00000001/art00006.

Wilkinson:2012:PAN

[Wil12] James H. Wilkinson. The Pilot ACE at the National Physical Laboratory. In Copeland [Cop12], pages 93–104. ISBN 0-19-960915-2 (paperback). LCCN ???? UK £14.99.

Wilkinson:2023:REA

[Wil23] J. H. (James Hardy) Wilkinson. Rounding Errors in Algebraic Processes, volume 89 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2023. ISBN 1-61197-751-7. xiii + 161 pp. LCCN QA275 .W64 2023. New foreword by N. J. Higham.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.